# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ 4.4.1 Discussion: thermal noise and the late time behaviour of active particles
For $D=0$, the terms in the square bracket reduce to $D_{e}\,x_{0}$, similar to passive particles; and this flux is due to the particles with instantaneous negative spin. For nonzero but small $D$, $e^{-k_{a}x_{0}}$ is still negligible and the flux gains a contribution that couples thermal noise and activity: $\frac{Dv}{2\,\omega}(m_{0}+1-\frac{\omega D}{v^{2}})=D(\,\frac{v}{\omega}s_{+}-\frac{D}{2\,v})$. If $s_{+}>0$ the change is proportional to $\frac{Dv}{\omega}$ which is positive and the particles with initial positive spin largely contribute to an additional flux at the origin. This is consistent with the picture that at short times the particles with initial positive spin move away from the origin, which makes them available in the system to contribute to the noise driven flux at large times. There is also a spin-independent reduction of the flux amounting to $-\frac{D^{2}}{2\,v}$, which is much smaller but gives the dominating correction for $s_{+}=0$. It pertains to the diffusive spread $\sim O(\frac{D}{v})$ over the persistent motion.
@@
[…]
which is positive at all times and undergoes a slow algebraic decay. It is tempting to relate this result to the simple fact that the particles with negative spin are more prone to go out of the absorbing boundary at $x=0$. However, quite remarkably, $Q^{\infty}(t)$ vanishes in the absence of thermal diffusion, implying that there are equal numbers of positive and negative spin particles in the system.
@@
The picture is considerably revised in the presence of even a small thermal diffusion. For very small $x$ the time to hit the absorbing boundary is small. But at such small times diffusion dominates over drift, and therefore it compels a larger fraction of the particles near $x=0$ to get absorbed. However the region near $x=0$ is majorly populated by the negative spin particles, implying their enhanced elimination from the system because of the diffusion induced absorption consistent with the flux reduction for $s_{+}=0~{}(m_{0}=-1)$ discussed after Eq. (70). This leads to an imbalance of the two species of particles. This also results in removing the discontinuity at the boundary through the formation of the kinetic boundary layer, which in this case is a depletion layer, of width $k_{a}^{-1}\sim D/v$ for small $D$. As a whole, diffusion has a complex and often nonnegligible effect on the behaviour of active particle dynamics. Notably, the bulk density profile in Eq. (66) deviates from the absorbing condition at $x=0$ by an amount proportional to $\frac{v^{2}}{\omega\,\sqrt{2\omega D_{e}}}$, i.e. the Milne length, which is precisely compensated by the boundary layer.
5 Universality in the large time distribution
The density profile for the absorbing boundary problem of underdamped passive dynamics resembles closely the distribution in Eq. (66), except that in the former case the boundary layer structure is richer and a boundary discontinuity is present. The solution for the underdamped case is complicated and originally given in the Laplace space [59]. However it is relatively simpler to extract the approximate large time behaviour, which is shown in the Appendix C. Note that the underdamped passive problem is exactly equivalent to the absorbing boundary problem of active Ornstein-Uhlenbeck process (AOUP) in the absence of translational noise and external forces, and so are the solutions. The striking resemblance of such disparate processes, one of which is a passive process, with the only common point being the presence of an additive exponentially correlated noise evokes a general form for the large time distribution in presence of an absorbing barrier,
@@ 5 Universality in the large time distribution
$+~{}{\rm Boundary~{}layer},$
where $u_{0}$ is the average initial velocity, $\tau_{p}$ is the noise correlation time or persistence time, and $l_{M}$ is the Milne length. Clearly, the distribution has three components, $P_{\rm tr}^{\infty}(x,t)=P_{\rm BM}+P_{\rm CS}+P_{\rm BL}$, where $P_{\rm BM}$ is the passive-like contribution with an effective diffusivity, $P_{\rm BL}$ is the dynamics dependent boundary layer, and $P_{\rm CS}$ is proposed as a universal scaling form emergent due to the coloured noise, $P_{\rm CS}(x,t;x_{0})=\frac{l_{M}}{D_{e}t}\,\psi(\frac{x}{\sqrt{D_{e}t}};\frac{x_{0}}{\sqrt{D_{e}t}})$ with,
@@
A nonzero Milne length implies that the particle far from the boundary sees the absorbing condition at $x=-l_{M}$ instead of $x=0$. However in Eq. (66) or Eq. (72), the quantity $P_{\rm tr}^{\infty}-P_{\rm BL}$ doesn’t exactly vanish at $-l_{M}$ and is off from zero by a tiny amount $\sim l_{M}^{2}t^{-3/2}~{}{\rm or}~{}l_{M}\,u_{0}\tau_{p}\,t^{-3/2}$. For RTP such terms occur from $O(L^{-2})$ corrections in the large-$L$ expansion of the spectrum and are omitted. Motivated from Eq. (C.2) for the underdamped passive case, we incorporate these and combine with $P_{\rm CS}$ and $P_{\rm BM}$ to write the distribution at large times $\sqrt{D_{e}t}\gg u_{0}\tau_{p},l_{M}$ in a particularly simple form,
@@
where $\pm$ is for $x\lessgtr x_{0}$ respectively. This particular form correctly takes into account the absorbing condition at $x=-l_{M}$, and is proposed as the universal distribution for the absorbing boundary problem for dynamics with colored noise with short-ranged correlations. Using simulations we have checked the validity of Eqs. (74) for nonthermal ABP as well, which is shown in the Appendix D.
@@ 6 Conclusion
Analytical results discussed in this paper for boundary driven noninteracting RTPs provide several new insights for nonequilibrium processes subject to an exponentially correlated noise. For zero boundary magnetisations, which is closest to the passive case, the steady state current satisfy the Fourier law, and in the large $L$ limit both the bulk density profile and current goes over to the usual diffusive behaviour. Notably, in contrast to the passive case, a kinetic boundary layer and finite bulk magnetisation emerges in the system, and the bulk density acquires important corrections equivalent to the substitutions: $x\rightarrow x+l_{M}$ and $L\rightarrow L+2\,l_{M}$, reminiscent of an effective absorbing conditions outside the boundaries. The scenario with nonzero boundary magnetisation is however starkly different in almost all aspects: The Fourier law is violated, the bulk density and the Milne length can no longer be related to an effective boundary condition, and thermal diffusion acquires a profound role. While nontrivial boundary features are generic to active processes, the complex interplay of the active motion and thermal noise leads to unexpected current reversal in the presence of diffusion; surprisingly, the current also shows a nonmonotonous behaviour as the diffusivity and activity parameters are tuned. Whether such current reversal and nonmonotonous transport are generic to boundary driven active systems will be worth studying.
In presence of finite boundary magnetisations a Seebeck-like effect is identified. In the equilibrium Seebeck effect, the particles carry charge and also energy, the latter being distributed according to the temperatures at the boundaries. It is plausible that if the particles carry more than one attributes relevant to the dynamics, as in the underdamped as well as the active cases, such effects might occur. An intriguing question is whether magnetisation, which doesn’t have any passive counterpart nor has an associated conserved flux, may still be regarded as a thermodynamic force.
@@
The striking similarity of the distribution in the absorbing boundary problem of such diverse systems points to a new universality in the physics induced by exponentially correlated noise. The unconstrained motions of these particles exhibit diffusive behaviour at large times; similar conclusion holds for reflecting boundaries [28]. The active and passive processes are however very different in the presence of interaction, and a set of model independent features for active single-file dynamics were found recently [60]. Contrary to both of the above, in the absorbing boundary problem the proposed universal distribution in Eq. (74) is new and different from the diffusive one, and it appears to hold for any exponentially correlated noise including the passive underdamped motion. We further expect that the steady state of the boundary driven problem in these cases will have a structure similar to Eq. (7), i.e. sum of a linear profile and boundary layers [56, 57]. We must however be cautious that such similarities of active and underdamped passive cases might be lost in presence of trapping and other forces, other noises, or interactions  [^4] For example the steady state of a trapped athermal AOUP is different from that of a trapped underdamped passive particle [61].. This is because the active motion considered here is overdamped which is different from the underdamped motion in general, and it would be worthwhile to look into the interplay of inertia and active noise.
Keeping the differences in mind, it seems plausible that the formation of a boundary layer in presence of fluxes at boundaries is generic for motions with coloured noise. For that matter the noise having continuous values might be important, specifically for the athermal case. The one dimensional athermal AOUP gives rise to boundary layers (along with a boundary discontinuity), while the athermal RTP results in a boundary discontinuity only which is smoothened to a boundary layer in presence of diffusion. For the RTP in a 1D box with reflecting boundaries a delta peak occurs at the walls that is converted to an exponential boundary layer for $D>0$ [13]. On the other hand, in 2-dimension where the orientation of RTP takes continuous values, a boundary layer is reported for reflecting boundaries even when thermal noise is absent [62], and we would expect a kinetic boundary layer for absorbing boundary. Studying absorbing boundary problems with other realisations of continuous valued coloured noise would be useful in this regard.
@@
In conclusion, boundary driven active dynamics provides an elementary physical mechanism for diverse nonequilibrium phenomena -- some of which are reported here for the first time, while the others were found previously in disparate settings. This work specifically elucidates the interplay of thermal noise with activity, highlights new features in boundary driven transport, and suggests a novel universality for the motion involving coloured noise. Whether these features survive when the noise correlation is non-exponential yet short ranged, as well as the qualitative changes introduced by interactions, will be of particular interest.
@@ Acknowledgements
AD acknowledges discussions with Soumyadip Banerjee.
Appendix A Exit probability of RTP at the left boundary
The exit probabilities $E_{\pm}(x)$ for an RTP is obtained by solving two coupled backward equations with boundary conditions, $E_{\pm}(0)=1,E_{\pm}(L)=0$ (see Eq. (51) of [13] with signs interchanged). The equations are identical to that satisfied by $P_{\mp}(s)$ in the steady state and for zero boundary magnetisations the density and magnetisation profiles can be written as,
$P(x)=P_{1}-\Delta P\,[E_{+}(x)+E_{-}(x)],~{}Q(x)=\Delta P\,[E_{+}(x)-E_{-}(x)],$ (A.1)
where $\Delta P=P_{1}-P_{0}\,$. This relation is used for simulating the steady state distributions. It however turns out that, as the tumble dynamics mixes the spin states, the correspondence between the exit probabilities and the steady state distributions are lost for general boundary conditions.

### v4

@@ 4.5.1 Discussion: thermal noise and the late time behaviour of active particles
For $D=0$, the terms in the square bracket reduce to $D_{e}\,x_{0}$, similar to passive particles; and this flux is due to the particles with instantaneous negative spin. For nonzero but small $D$, $e^{-k_{a}x_{0}}$ is still negligible and the flux gains a contribution that couples thermal noise and activity: $\frac{Dv}{2\,\omega}(m_{0}+\frac{v}{\sqrt{2\,\omega\,D_{e}}})\approx\frac{Dv}{2\,\omega}(m_{0}+1-\frac{\omega D}{v^{2}})=D(\,\frac{v}{\omega}s_{+}-\frac{D}{2\,v})$. If $s_{+}>0$ the change is proportional to $\frac{Dv}{\omega}$ which is positive and the particles with initial positive spin largely contribute to an additional flux at the origin. This is consistent with the picture that at short times the particles with initial positive spin move away from the origin, which makes them available in the system to contribute to the noise driven flux at large times. There is also a spin-independent reduction of the flux amounting to $\frac{D^{2}}{2\,v}$, which is much smaller but gives the dominating correction for $s_{+}=0$. It pertains to the diffusive spread $\sim O(\frac{D}{v})$ over the persistent motion.
@@
$Q^{\infty}(t)=\int_{0}^{\infty}Q_{\rm tr}^{\infty}(x,t)\,dx=\frac{v\,D}{\omega\sqrt{2\omega D_{e}}}\,\frac{2\pi}{(4\pi D_{e}t)^{3/2}}\,x_{0}\,e^{-\frac{x_{0}^{2}}{4D_{e}t}},$ (65)
which is positive and undergoes a slow algebraic decay with time. It is tempting to relate this result to the simple fact that the particles with negative spin are more prone to go out of the absorbing boundary at $x=0$. However, quite remarkably, $Q^{\infty}(t)$ vanishes in the absence of thermal diffusion, implying that there are equal numbers of positive and negative spin particles in the system.
@@
The picture is considerably revised in the presence of even a small thermal diffusion. For very small $x$ the time to hit the absorbing boundary is small. But at such small times diffusion dominates over drift, and therefore it compels a larger fraction of the particles near $x=0$ to get absorbed. However the region near $x=0$ is majorly populated by the negative spin particles, implying their enhanced elimination from the system because of the diffusion induced absorption consistent with the flux reduction for $s_{+}=0~(m_{0}=-1)$ discussed after Eq. (64). This leads to an imbalance of the particles with different orientations. This also results in removing the discontinuity at the boundary through the formation of the kinetic boundary layer, which in this case is a depletion layer, of width $k_{a}^{-1}\sim D/v$ for small $D$. As a whole, diffusion has a complex and often nonnegligible effect on the behaviour of active particle dynamics. Notably, the bulk density which is referred to as the ‘scaling’ part of the profile in Eq. (60), when extended to $x=0$, deviates from the absorbing condition by an amount proportional to $l_{M}=\frac{v^{2}}{\omega\,\sqrt{2\omega D_{e}}}$. This deviation is precisely compensated by the boundary layer.
5 Proposed universality in the large time distribution
A natural question is whether the new features in the distribution is specific to RTPs or is more general to active processes. We propose that not only the new features are general, but in fact the large time distribution in Eq. (60) excepting the boundary layer is universal for particle motion subject to exponentially correlated noise.
Let us consider for example the absorbing boundary problem of underdamped passive dynamics. The solution for the underdamped case is complicated and originally given in the Laplace space [60]. However it is relatively simpler to extract the approximate large time behaviour, which is shown in the Appendix E. Quite remarkably the density profile in that case is precisely of the form given in Eq. (60), except with a richer boundary layer structure (Eq. (E.3)) and a boundary discontinuity. Note that the underdamped passive problem is exactly equivalent to the absorbing boundary problem of active Ornstein-Uhlenbeck process (AOUP) in the absence of translational noise and external forces, and so are the solutions. The striking resemblance of such disparate processes, one of which is a passive process, with the only common point being the presence of an additive exponentially correlated noise evokes a general form for the large time distribution in presence of an absorbing barrier,
@@ 5 Proposed universality in the large time distribution
$+~P_{\rm BL},$
where $u_{0}$ is the average initial velocity, $\tau_{p}$ is the noise correlation time or persistence time, $l_{M}$ is the Milne length, and $P_{\rm BL}$, which is nonuniversal and generically falls off exponentially with distance, represents the boundary layers. Clearly, the distribution has three components, $P_{\rm tr}^{\infty}(x,t)=P_{\rm BM}+P_{\rm CS}+P_{\rm BL}$. Here $P_{\rm BM}$ is the passive-like contribution with an effective diffusivity, $P_{\rm BM}=\frac{1}{\sqrt{4\pi D_{e}t}}\left\{e^{-\frac{(x-x_{0})^{2}}{4D_{e}t}}-e^{-\frac{(x+x_{0})^{2}}{4D_{e}t}}\right\}$ and $P_{\rm CS}$ is proposed as a universal scaling function due to the coloured noise, $P_{\rm CS}(x,t;x_{0})=\frac{l_{M}}{D_{e}t}\,\psi(\frac{x}{\sqrt{D_{e}t}};\frac{x_{0}}{\sqrt{D_{e}t}})$,
@@
At this stage we discuss whether $l_{M}$ indeed is the Milne length, which would imply that the particle far from the boundary sees the absorbing condition at $x=-l_{M}$ instead of $x=0$, as in the steady state with $Q_{0}=Q_{1}=0$ (section 3.1). In Eq. (60) or Eq. (66), the quantity $P_{\rm tr}^{\infty}-P_{\rm BL}$ doesn’t exactly vanish at $l_{M}$ and is off from zero by a tiny amount $\sim l_{M}^{2}t^{-3/2}~{\rm or}~l_{M}\,u_{0}\tau_{p}\,t^{-3/2}$. For RTP such terms occur from $O(L^{-2})$ corrections in the large-$L$ expansion of the spectrum and were omitted. Motivated from Eq. (E.2) for the underdamped passive case, we incorporate these in Eq. (66) to write the distribution at large times $\sqrt{D_{e}t}\gg u_{0}\tau_{p},l_{M}$ in a particularly simple form,
@@
where $\pm$ is for $x\lessgtr x_{0}$ respectively. This particular form correctly takes into account the absorbing condition at $x=-l_{M}$, and is proposed as the universal distribution for the absorbing boundary problem for dynamics with colored noise with short-range correlations. Using simulations we have checked the validity of Eq. (68) for nonthermal ABP as well, which is shown in the Appendix F.
@@ 6 Conclusion
Analytical results discussed in this paper for noninteracting free RTPs driven by active reservoirs provide several new insights for nonequilibrium processes subject to an exponentially correlated noise. For zero boundary magnetisations, which is closest to the passive case, the steady state current satisfy the Fick’s law, and in the large $L$ limit both the bulk density profile and current show to the usual diffusive behaviour. Notably, in contrast to the passive case, a kinetic boundary layer and finite bulk magnetisation emerges in the system. The bulk density acquires important corrections equivalent to the substitutions: $x\rightarrow x+l_{M}$ and $L\rightarrow L+2\,l_{M}$ with $l_{M}$ the Milne length, reminiscent of effective absorbing conditions outside the boundaries. The scenario with nonzero boundary magnetisation is however starkly different in almost all aspects: The Fick’s law is modified and a nontrivial particle current and bulk magnetisation are induced solely due to the boundary magnetisations. Generating steady currents by ‘pushes’ localised at the boundaries, unlike thermodynamic potential having nonzero gradients in the bulk, is not obvious in such noninteracting unbiased dynamics. In this case the bulk density profile can no longer be related to an effective boundary condition, and thermal diffusion acquires a profound role. While nontrivial boundary features are generic to active processes, the complex interplay of the active motion and thermal noise can lead to unexpected current reversal which in turn would cause a change in the nature of the distribution. Surprisingly, the current also shows a nonmonotonous behaviour as the diffusivity and activity parameters are tuned. Whether such current reversal and nonmonotonous transport are generic to boundary driven active systems will be worth studying. It may also provide avenues to manoeuvre the system without affecting the boundary conditions.
In presence of boundary magnetisations a new transport behaviour emerges, which manifests only at a global scale and loosely resembles a Seebeck-like effect. A key question is the nature of magnetisation which plays a dual role. The boundary magnetisation, although does not have an associated conserved flux, induce a particle current with magnitude $O(L^{-1})$ similar to a potential. On the other hand the small and constant bulk magnetisation teams up with the density gradient in driving the current, thus resembling a force. Whether magnetisation and such quantities can act as novel nonequilibrium thermodynamic attributes might be an intriguing question. Devising a boundary magnetisation regulator and simulating particle-boundary interaction remain open, and some of the possibilities outlined in the Appendix A will be explored in future.
@@
The striking similarity of the distribution in the absorbing boundary problem of such diverse systems points to a new universality in the physics induced by exponentially correlated noise. The unconstrained motions of these particles exhibit diffusive behaviour at large times; similar conclusion holds with reflecting boundaries [28]. The active and passive processes are however very different in the presence of interaction, and a set of model independent features for active single-file dynamics were found recently [61]. Contrary to both of the above, in the absorbing boundary problem the proposed universal distribution in Eq. (68) is new and different from the diffusive one, and it is expected to hold for any exponentially correlated noise including the passive underdamped motion. We further expect that the steady state of the boundary driven problem in these cases will have a structure similar to Eq. (7), i.e. sum of a linear profile and boundary layers [57, 58]. We must however be cautious that such similarities of active and underdamped passive cases might be lost in presence of trapping and other forces, other noises, or interactions  [^9] For example the steady state of a trapped athermal AOUP is different from that of a trapped underdamped passive particle [62].. This is because the active motion considered here is overdamped which is different from the underdamped motion in general, and it would be worthwhile to look into the interplay of inertia and active noise.
Keeping the differences in mind, it seems plausible that the formation of a boundary layer in presence of fluxes at boundaries is generic for motions with coloured noise. For that matter the noise having continuous values might be important, specifically for the athermal case. The one dimensional athermal AOUP gives rise to boundary layers (along with a boundary discontinuity), while the athermal RTP results in a boundary discontinuity only which is smoothened to a boundary layer in presence of diffusion. For the RTP in a 1D box with reflecting boundaries a delta peak occurs at the walls that is converted to an exponential boundary layer for $D>0$ [13]. On the other hand, in 2-dimension where the orientation of RTP takes continuous values, a boundary layer is reported for reflecting boundaries even when thermal noise is absent [63], and we would expect a kinetic boundary layer for absorbing boundary. Studying absorbing boundary problems with other variations of continuous valued coloured noise would be useful in this regard.
@@
In conclusion, boundary driven active dynamics provides an elementary and unifying physical ground for diverse nonequilibrium phenomena - some of which are reported here for the first time, while the others were found previously in disparate settings. This work specifically elucidates the interplay of thermal noise with activity, highlights new features in boundary driven transport, and suggests a novel universality for the motion involving coloured noise. Whether these features survive when the noise correlation is non-exponential yet short ranged, as well as the qualitative changes introduced by interactions, will be of particular interest.
@@ Acknowledgements
AD acknowledges discussions with Soumyadip Banerjee. The authors thank the reviewers for their constructive comments which helped to improve the manuscript.
Appendix A Comments on the boundary magnetisation control
Realising the boundary magnetisations in specific systems is important for observing the behaviour reported in the paper, which requires modelling the reservoir and the reservoir-particle interaction at the boundaries. Active particles are known to have nontrivial boundary features in general, and it is expected that the boundary magnetisation is typically nonzero in any realistic scenario. Here we qualitatively outline few setups for its control, which is speculative at this stage. Overdamped dynamics and the presence of thermal noise is assumed.
An imbalance of concentration of particles with different orientations may be achieved by modelling the reservoir as RTPs confined in a rectangular box. The box width must be much larger than that of the channel (Figure 1), while the channel width should be significantly larger than the particle dimensions and a dilute limit to be realised to minimise interactions. The differently oriented particles accumulate at different sides of the reservoir box, e.g. at the right end of the reservoirs particles with predominantly positive spins accumulate. This provide a nonzero magnetisation at reservoir-channel junction. Following [13] we expect that for box length large compared to other relevant length scales related to particle size and motility, the magnetisation and density at its edges behave differently with changing diffusivity. This can give a handle to externally control different boundary conditions in a certain range. Introducing a partially permeable sticky membrane [64] in the reservoir or imposing a local bias near the junction might help to tune these quantities further. These might also be useful for making an ‘active-couple’ mentioned before.
To facilitate simulations we may view the problem as follows. A density $P_{0}=P_{0}^{+}+P_{0}^{-}$ and magnetisation $Q_{0}=P_{0}^{+}-P_{0}^{-}$ can be rephrased as the conditional probability $s_{0}^{\pm}=\frac{P_{0}^{\pm}}{P_{0}}$ of finding the particle with spin $\pm 1$ respectively, at the given point; this also induces a net local velocity $u=v\,(s_{0}^{+}-s_{0}^{-})=v\,\frac{Q_{0}}{P_{0}}$ to the particle. One may think of enforcing it, e.g. by manufacturing preferential ‘kicks’ or manipulating the tumble dynamics in the reservoir or at the boundaries. Similar ideas were employed to study RTP in a box with reflecting walls where the tumble occurs only at the boundaries [65]. Devising an appropriate boundary interaction in the presence of bulk tumbling, boundary fluxes and thermal noise is however challenging. Indeed, modelling particle exchange with reservoirs for systems defined on continuum and acted on by thermal noise is a formidable task [32], which largely prevents simulating such processes. Implementing the problem on a lattice maybe better suited for simulations at present.
On the other hand, in the zero current (or ‘active-couple’) case we can avoid the particle exchange with reservoir and work in a closed setting. Now the densities $P^{\pm}$ at the boundaries are not known except their difference in the steady state, $Q_{0,1}=P^{+}_{0,1}-P^{-}_{0,1}$. The task is to model a suitable particle-boundary interaction. Since this is a closed system and the dynamics is symmetric, $1=\int P(x)\,dx=\frac{P_{0}+P_{1}}{2}L$ in the steady state. Using this along with the boundary magnetisations, we can infer the steady state boundary values. Then we can enforce time independent probabilities $s_{0,1}^{\pm}=\frac{P^{\pm}_{0,1}}{P_{0,1}}\rvert_{\rm ss}$ of particle orientations at the boundaries. Simply put, in simulations we can specify $s_{0,1}^{\pm}$ and impose reflecting boundary conditions. Finding the distributions theoretically in this setup might be interesting.
Appendix B Exit probability of RTP at the left boundary
The exit probabilities $E_{\pm}(x)$ at the left boundary for an RTP initially at $x$ with spin $\pm 1$ are obtained by solving two coupled backward equations with boundary conditions, $E_{\pm}(0)=1,E_{\pm}(L)=0$ (see Eq. (51) of [13] with signs interchanged). The equations are identical to that satisfied by $P_{\mp}(x)$ in the steady state and for zero boundary magnetisations the density and magnetisation profiles can be written as,
$P(x)=P_{1}-\Delta P\,[E_{+}(x)+E_{-}(x)],~Q(x)=\Delta P\,[E_{+}(x)-E_{-}(x)],$ (B.1)
where $\Delta P=P_{1}-P_{0}\,$. After simulating the exit probability, this relation is used for finding the steady state distributions with boundary conditions $Q_{0}=Q_{1}=0$. It however turns out that, as the tumble dynamics mixes the spin states, the correspondence between the exit probabilities and the steady state distributions are lost for general boundary conditions.